OpenStudy (anonymous):

Q-12 Find the angle of elevation of the sun when a 6m high pole makes shadow of length 2squrt3m on the horizontal plane. a. 300 b. 450 c. 600 d. 900

OpenStudy (amistre64):

tan-1(6/2sqrt(3))

OpenStudy (anonymous):

you should write "degrees"

OpenStudy (amistre64):

c 600 degrees :)

OpenStudy (anonymous):

I mean the asker

OpenStudy (anonymous):

"60 degrees" , not 600

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

can u plz draw

OpenStudy (amistre64):

no, im sure that you are capable enough to pick up your own pencil and paper and draw a picture. you really dont need anyone to hold your hand in doing that

OpenStudy (anonymous):

its fairly simple lol

OpenStudy (amistre64):

lol.... cute pic ;)

OpenStudy (amistre64):

but its not to scale.... draw it again ..

OpenStudy (anonymous):

tan-1(6/2squrt3) is it a formula?

OpenStudy (amistre64):

no, it is an operation. just like + or / is an operation

OpenStudy (amistre64):

tan-1(opp/adj) would be a formula

OpenStudy (anonymous):

u mean perp/base

OpenStudy (amistre64):

tan(a) = m/n tan-1(tan(a)) = tan-1(m/n) a = tan-1(m/n)

OpenStudy (amistre64):

perp/base is a suitable interpretation for this problem yes

OpenStudy (anonymous):

it is tan inverse?

OpenStudy (amistre64):

it is :)

OpenStudy (anonymous):

ok ok ok thanks amitre 4 ur humble response

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

in the end the answer wil b squrt3 yes?

OpenStudy (amistre64):

in the end; the answer is 60 degrees

OpenStudy (anonymous):

u dnt know what ur doing :|

OpenStudy (amistre64):

me? lol .... you might be right

OpenStudy (anonymous):

remember u are meant to learn from this site, not just get answers and

OpenStudy (anonymous):

no, was talking to the asker

OpenStudy (anonymous):

i mean that then tan-1 (squrt3)=60deg

OpenStudy (anonymous):

I can tell some people on this site are only after answers. Especially when they say things like "is that the final answer" , because that means that you are sort of only after the final answer and you didnt understand the reasoning and working out.

OpenStudy (amistre64):

close .... \(\tan^{-1}(6/(2\sqrt{3}))=60^o\)

OpenStudy (anonymous):

ok amister

OpenStudy (amistre64):

.... and yes; tan-1(sqrt(3)) does = 60 degrees; but i was thinking in the context of the problem posted